Find the inverse of f(x) = 2/5-x
Answer by stanbon(75887)

Find the inverse of f(x) = 2/5-x
-----
Do you mean y = (2/5)-x or do you mean y = 2/(5-x)
-------
To find the inverse::
1. Interchange x and y
x = (2/5)-y
2. Solve for "y"::
y = (2/5)-x
-------
Or, if you mean y = 2/(5-x)
1. Interchange x and y
x = 2/(5-y)
2. Solve for "y"::
5-y = 2/x
y = (2/x)+5
----------------
Cheers,
Stan H.
---------------